40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 29 01:52:22 2017, gautam, Update, SUS, Test mass actuator calibration ALS_comparison.pdf
    Reply  Tue Aug 29 15:04:59 2017, gautam, Update, SUS, ETMY Oplev PIT loop gain changed ETMY_OLPIT.pdfGTR_comparison.pdf
    Reply  Wed Oct 18 17:34:09 2017, gautam, Update, SUS, ASDC 
Message ID: 13265     Entry time: Tue Aug 29 01:52:22 2017     Reply to this: 13267   13392
Author: gautam 
Type: Update 
Category: SUS 
Subject: Test mass actuator calibration 

[ericq, gautam]

Tonight, we decided to double-check the POX counts-to-meters conversion.

It is unclear when this was last done, and since I modified the coil driver electronics for the ITMs and BS recently, I figured it would be useful to get this calibration done. The primary motivation was to see if we could resolve the discrepancy between the current ALS noise (using POX as a sensor) compared to the Izumi et. al. plot.

Because we are planning to change the coil driver electronics further soon anyways, we decided to do the calibration at a single frequency for tonight. For future reference, the extension of this method to calibrate the actuator over a wider range of frequencies is here. The procedure followed, and the relevant numbers from tonight, are as follows.

Procedure:

  1. Set dark offsets on all DCPDs and LSC PDs.
  2. Look at the free swinging Michelson signal on ASDC.
    • For tonights test, ASDC was derived from the AS55 photodiode.
    • The AS110 photodiode actually has more light on it, but we think that the ADC that the DCPD board is interfaced to is running on 0-2V rather than 0-10V, as the signal seemed to saturate around 2000 counts. It is unclear whether the actual photodiode is saturating, to be investigated.
    • So we decided to use ASDC from AS55 photodiode with 15dB whitening gain.
    • There is also some issue with the whitening filter (not whitening gain) on ASDC - engaging the whitening shifts the DC offset. This has to be investigated while we get stuck into the LSC electronics.
  3. Look at the peak-to-peak swing of ASDC. Use algebraic expression for reflected power from Michelson interferometer to calibrate the ASDC slope at Michelson half-fringe. For the test tonight, ASDC_max = 1026 counts, ASDC_min = 2 counts.
  4. Lock the Michelson at half-fringe, with ASDC as the error signal.
    • Zero out the MICH elements in the RFPD input matrix.
    • Set the matrix element from ASDC to MICH in the DCPD LSC input matrix to 1.
    • The servo gain used was +0.005 on the MICH_A servo path.
    • A low-frequency boost was turned on.
  5. Use the sensing matrix infrastructure to drive a line in the optic of interest.
    • Tonight, we looked at ITMX and ITMY.
    • The line was driven at 311.1Hz, and the amplitude was 300 counts.
    • Download 60secs of ASDC data, demodulate at the driven frequency to find the peak height in counts, and using the slope of ASDC (in cts/m) at the Michelson half-fringe, calculate the actuator gain in m/cts.
    • ITMY: 2.55e-9 / f^2 m/count
    • ITMX: 2.65e-9 / f^2 m/count
    • These numbers kind of make sense - the previous numbers were ~5nm/f^2 /ct, but I removed an analog gain of x3 in this path. Presumably there has been some change in the N/A conversion factor - perhaps because of a change in the interaction between the optics' face magnets and the static magnetic field in the OSEMs?
  6. Lock the arms with POX/POY, and drive the newly calibrated ITMs.
    • So we know how many meters we are driving the ITMs by.
    • Looking at POX/POY, we can calibrate these into meters/count.
    • Both POX and POY were whitened.
    • POX whitening gain = +30dB, POY whitening gain = +18dB.
    • ITMX and ITMY were driven at 311.1Hz, with amplitude = 2counts.
    • Download 60 secs of data, demodulate at the drive frequency to find the peak height, and use the known ITM actuator gains to calibrate POX and POY.
    • POX: 7.34e-13 m / count (approx. 5 times less than the number in the Foton filter bank in the C1:CAL-CINV model).
    • POY: 1.325e-13 m / count
    • We did not optimize the demod phases for POX/POY tonight. 

Once these calibrations were updated, we decided to control the arms with ALS, and look at the POX spectrum. Y-arm ALS wasn't so stellar tonight, especially at low frequencies. I can see the GTRY spot moving on the CCD monitor, so something is wonky. To be investigated. But the X arm ALS noise looked pretty good.

Seems like updating the calibration did the job; see the attached comparison plot.

Attachment 1: ALS_comparison.pdf  177 kB  Uploaded Tue Aug 29 03:25:56 2017  | Hide | Hide all
ALS_comparison.pdf
ELOG V3.1.3-